Twenty-four karat gold is considered pure. If a coin is made of 22 karat gold, what is its purity as a fraction and as a decimal rounded to the nearest thousandth?
Answer by TimothyLamb(4379)

---
purity fraction:
22/24 = 11/12
---
purity decimal:
11/12 = 0.917
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php